# Feferman’s completeness theorem

### v1

@@ 4. A theorem of Ash and Knight in $\mathsf{ACA_{0}}$
The goal of this section is to prove a version of a theorem of Ash and Knight [ash1990] from computable structure theory in $\mathsf{ACA_{0}}$. In its original formulation, this theorem connects membership in hyperarithmetical sets to the well orderedness of certain linear orders. Before we state and prove our version of this theorem we need to make some definitions.
We are now ready to state the version of Ash and Knight’s theorem we have in mind.
@@
In order to prove Theorem 10 we will need a few more lemmas. The first is the relativized version of Shoenfield’s classical limit lemma. See [soare1999] for the classical proof of this theorem and more background on computability theory.
@@
A linear order $\mathcal{L}=(L,\preceq^{\mathcal{L}})$ is $\Delta^{0}_{n}(X)$ if $\preceq^{\mathcal{L}}$ is $\Delta^{0}_{n}(X)$. Notice that this implies that the universe $L$ is also $\Delta^{0}_{n}(X)$, since by reflexivity $x\in L$ iff $x\preceq^{\mathcal{L}}x$. If $\mathcal{L}$ is $\Delta^{0}_{1}(X)$, then we say that $\mathcal{L}$ is $X$-computable. We will denote linear orders by calligraphic letters and their universes $L$ by the corresponding capital letters. A sequence of linear orders $(\mathcal{A}_{i})_{i\in\omega}$ is uniformly $\Delta^{0}_{n}(X)$ if there is a $\Delta^{0}_{n}(X)$ set $S$ such that its $i^{\text{th}}$ column is equal to $\preceq^{\mathcal{A}_{i}}$, i.e.,
@@
From this one can easily extract a pair of $\Pi^{0}_{2}$ formulas which given parameter $X$ define the membership relation of $Y$, respectively, its negation. Thus $Y$ is $\Delta^{0}_{2}(X)$. ∎
From Lemma 11 we can derive a limit lemma for linear orders. Recall that even if $\mathcal{L}$ is $\Delta^{0}_{2}(X)$, we assume that $L$ is $X$-computable.
We say that a sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$ is locally coherent if $\preceq^{\mathcal{A}_{i}}\upharpoonright(A_{i}\cap A_{i+1})=\preceq^{\mathcal{A}_{i+1}}\upharpoonright(A_{i}\cap A_{i+1})$, for each $i$. We will be interested in the situation when a linear order $\mathcal{L}$ is the limit of a locally coherent sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$. In this case we think about $\mathcal{L}$ as the order obtained by the following procedure. We start with the order $\mathcal{A}_{0}$. At each stage $i$ we switch from $\mathcal{A}_{i}$ to $\mathcal{A}_{i+1}$ by removing all the elements $x\in A_{i}\setminus A_{i+1}$ and adding all the elements in $x\in A_{i+1}\setminus A_{i}$. The order $\mathcal{L}$ consists of all the elements $x$ that are present in some $\mathcal{A}_{i}$ and aren’t removed in any $\mathcal{A}_{j}$, for $j>i$. Whether $x\preceq^{\mathcal{L}}y$ can be checked in any $\mathcal{A}_{i}$ such that both $x,y$ aren’t removed at any stage $\geq i$.
@@
Note that for a linear order without successor relation $\mathcal{L}=\lim\mathcal{A}_{i}$ and successive $x,y\in\mathcal{L}$, new elements $z$ might appear and disappear between $x$ and $y$ inside $\mathcal{A}_{i}$’s for indefinitely large $i$. However, for a linear order with successor relation $\mathcal{L}=\lim\mathcal{A}_{i}$ and neighboring elements $x,y\in\mathcal{L}$, for large enough stages (after the point when both $x$ and $y$ will no longer be removed) we could guarantee that no more new elements will appear in $\mathcal{A}_{i}$’s between $x$ and $y$.
@@
Fix a set $X$. A linear order $\mathcal{L}$ is $\Delta^{0}_{2}(X)$ if and only if there is an $X$-computable locally coherent sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$ with limit $\mathcal{L}$, i.e.,
$x\in L\leftrightarrow\exists i(\forall j>i)x\in A_{j}\text{ and }x\preceq^{\mathcal{L}}y\leftrightarrow\exists i(\forall j>i)x\preceq^{\mathcal{A}_{j}}y.$
@@
()($\Rightarrow$)( ⇒ ). From Lemma 11 we get a sequence $(B_{i})_{i\in\omega}$ of finite sets with limit $\preceq^{\mathcal{L}}$. We may assume that no $B_{i}$ contains elements greater than $i$. We define a locally coherent sequence $(\mathcal{A}_{i})_{i\in\omega}$ as follows. Let $\mathcal{A}_{0}$ be the empty order. Assume we have defined $\mathcal{A}_{i}$. To define $\mathcal{A}_{i+1}$ first let $C_{i+1}$ be the largest subset of $\{x:\langle x,x\rangle\in B_{i+1}\}$ linearly ordered by $B_{i+1}$. Then let $D_{i+1}$ be the largest subset of $C_{i+1}\cap A_{i}$ such that $B_{i+1}\upharpoonright D_{i+1}=\preceq^{\mathcal{A}_{i}}\upharpoonright D_{i+1}$ and define $\mathcal{A}_{i+1}$ by
$A_{i+1}=D_{i+1}\cup(C_{i+1}\setminus A_{i})\text{ and }\preceq^{\mathcal{A}_{i+1}}=B_{i+1}\upharpoonright A_{i+1}.$
This sequence is clearly uniformly $X$-computable. To see that it has limit $\mathcal{L}$, consider $x,y\in L$ such that $x\preceq^{\mathcal{L}}y$. There exists $i$ such that for all $j>i$ and $z,w\leq\max(x,y)$ we have $\langle z,w\rangle\in B_{j}\iff z\preceq^{\mathcal{L}}w$. By definition we have that $x\preceq^{\mathcal{A}_{i+2}}y$.
@@
Fix a set $X$. A linear order $\mathcal{L}$ with successor relation $S^{\mathcal{L}}$ is $\Delta^{0}_{2}(X)$ if and only if there is an $X$-computable locally coherent sequence of finite linear orders with successor relation $(\mathcal{A}_{i})_{i\in\omega}$ with limit $\mathcal{L}$.
We are now ready to prove two lemmas crucial for our proof of Theorem 10. They give a sufficient condition for the existence of computable linear orderings of order type $\omega\cdot\mathcal{L}$ for any linear ordering $\mathcal{L}$.
@@
Fix $X$ and let $\mathcal{L}$ be a $\Delta^{0}_{2}(X)$ linear order with a least element. Then there is an $X$-computable linear order of order type $\omega\cdot\mathcal{L}$ with an $X$-computable successor relation and 00 as its least element.
@@
Assume without loss of generality that 00 is the least element in $\mathcal{L}$. As $\mathcal{L}$ is $\Delta^{0}_{2}(X)$, by Corollary 12 there is a uniformly $X$-computable sequence of linear orders $(\mathcal{A}_{i})_{i\in\omega}$ with limit $\mathcal{L}$. We may furthermore assume that $\mathcal{A}_{i}$ does not contain elements greater than $i$ and that each $\mathcal{A}_{i}$ has the correct information about 00.
@@
We define $\mathcal{L}^{\prime}$ of order type $\omega\cdot\mathcal{L}$ as the union $\bigcup\limits_{s\in\mathbb{N}}\mathcal{L}_{s}^{\prime}$ of an $X$-computable sequence of finite linear orders $\mathcal{L}_{0}^{\prime}\subseteq\mathcal{L}_{1}^{\prime}\subseteq\ldots$. Together with $\mathcal{L}_{s}^{\prime}$ we define an $X$-computable sequence of order-preserving surjections $f_{s}\colon\mathcal{L}_{s}^{\prime}\to\mathcal{A}_{s}$.
We put $\mathcal{L}_{0}^{\prime}=\mathcal{A}_{0}$ and let $f_{0}$ be the identity function. Assume we have defined $\mathcal{L}_{s}^{\prime}$ and $f_{s}$. For every $x\in\mathcal{L}_{s}^{\prime}$, let $f_{s+1}(x)=f_{s}(x)$ if $f_{s}(x)\in\mathcal{A}_{s+1}$ and otherwise let $f_{s+1}(x)$ be the greatest element to the left of $f_{s}(x)$ in $\mathcal{A}_{s+1}\cap\mathcal{A}_{s}$, i.e.,
$f_{s+1}(x)=\max_{\mathcal{\preceq}^{\mathcal{A}_{s}}}\{a\in\mathcal{A}_{s+1}\cap\mathcal{A}_{s}:a\preceq^{\mathcal{A}_{s}}f_{s}(x)\}.$
Notice that $f_{s+1}$ is an order-preserving surjection from $\mathcal{L}_{s+1}$ onto $\mathcal{A}_{s+1}$.
At last, for each $a\in A_{s+1}$ we add to $\mathcal{L}_{s+1}$ an additional element $x_{a}$ such that $x_{a}$ is $\mathcal{L}_{s+1}$-greater than all previously present $y\in\mathcal{L}_{s+1}$ with $f_{s+1}(y)\preceq^{\mathcal{A}_{s+1}}a$ and $x_{a}$ is $\mathcal{L}_{s+1}$-smaller than all all previously present $y\in\mathcal{L}_{s+1}$ with $f_{s+1}(y)\succ^{\mathcal{A}_{s+1}}a$. We put $f_{s+1}(x_{a})=a$.
Verification: Clearly, the described procedure gives us computable sequences of $\mathcal{L}_{s}^{\prime}$ and $f_{s}$. Notice that for $a\in\bigcap\limits_{r\geq s}\mathcal{A}_{r}$ (that is $a$ that stabilizes in $\mathcal{A}_{r}$’s starting at the stage $r=s$) we have $f^{-1}_{s}(a)\subseteq f^{-1}_{s+1}(a)\subseteq f^{-1}_{s+2}(a)\subseteq\ldots$. Moreover, for $r\geq s$, the set $f^{-1}_{r+1}(a)\setminus f^{-1}_{r}(a)$ is a non-empty set such that all it’s elements are to the right (in $\mathcal{L}_{r+1}$) of the elements of $f^{-1}_{r}(a)$. We put $f\colon\mathcal{L}^{\prime}\to\mathcal{L}$ to be the function that maps $x$ to $f_{s}(x)$, where $s$ is the least number such that $x\in\mathcal{L}^{\prime}_{s}$ and $f_{s}(x)\in\bigcap\limits_{r\geq s}\mathcal{A}_{r}$. From the construction it is clear that $f$ is an order-preserving function and that for any $a\in\mathcal{L}$ the order type of $f^{-1}(a)$ inside $\mathcal{L}^{\prime}$ is $\omega$. Hence the order type of $\mathcal{L}^{\prime}$ is $\omega\cdot\mathcal{L}$. Notice that if $x\in\mathcal{L}_{s}^{\prime}$, then its successor in $\mathcal{L}^{\prime}$ is guaranteed to be in $\mathcal{L}_{s+1}^{\prime}$. This allows us to define an $X$-computable successor function on $\mathcal{L}^{\prime}$.
@@
Fix $X$ and let $\mathcal{L}$ be a $\Delta^{0}_{2}(X)$ linear order of order type $\omega\cdot\mathcal{K}$ for some linear order $\mathcal{K}$. Further assume that $\mathcal{L}$ has 00 as its least element and that the successor relation in $\mathcal{L}$ is $\Delta^{0}_{2}(X)$. Then there is an $X$-computable linear order isomorphic to $\mathcal{L}$ with 00 as its least element.
@@
As $\mathcal{L}$ is $\Delta^{0}_{2}(X)$, by Corollary 13 there is a uniformly $X$-computable locally coherent sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$ with partial successor relation and $\mathcal{L}$ as its limit. Furthermore we can assume without loss of generality that 00 is the least element of all $\mathcal{A}_{i}$.
@@
Verification: Clearly $\mathcal{L}^{\prime}$ is $X$-computable. It remains to show that $\mathcal{L}^{\prime}\cong\mathcal{L}$. First of all notice that the order $\mathcal{L}^{\prime}$ splits into the blocks $B_{a}$, for $a\in\mathcal{L}$, where $B_{a}=\bigcup\limits_{s\geq s_{a}}f_{s}^{-1}(a)$ and $s_{a}$ is the first stage such that $a\in\bigcap\limits_{s\geq s_{a}}A_{s}$. In order to finish the proof it is enough to show that all $B_{a}$ are finite and non-empty. For a given $a\in\mathcal{L}$ consider its successor $b$. In all the orders $\mathcal{A}_{s}$, for $s\geq\max(s_{a},s_{b})$, the element $b$ is the successor of $a$ according to $S^{\mathcal{A}_{s}}$. Thus no elements will appear in between $a$ and $b$ in the orders $\mathcal{A}_{s}$. Therefore no new elements will be added to $B_{a}$ after the stage $\max(s_{a},s_{b})$ and hence $B_{a}$ is finite. The block $B_{a}$ is non-empty since by construction $f^{-1}_{s_{a}}(a)=\{x_{a}\}$.∎
Combining Lemmas 14 and 15 we get a formal version of a special case of a theorem due to Ash [ash1991] as stated in [ash2000, Theorem 9.11]: If a linear order $\mathcal{L}$ has an isomorphic copy that is $\Delta^{0}_{3}$, then $\omega\cdot\mathcal{L}$ has a computable copy. The proof of this in two steps, first applying Lemma 14 with $X=\emptyset^{\prime}$ and then using Lemma 15, was communicated to us by Andrey Frolov.
@@
We can now build a uniformly $\Delta_{2n}$ sequence of linear orders $(\mathcal{C}_{i})_{i\in\omega}$ such that if $x\in X$, then $\mathcal{C}_{i}$ is the linear order only containing the element 00 and if $x\not\in X$, then $\mathcal{C}_{i}$ has order type $1+\eta$ where the first element is 00. The construction is again in stages where $\mathcal{C}_{i,s}$ contains only one element 00 if for all $y<s$ $\langle x,y\rangle\in Y$ and otherwise we build a copy of $\eta$ after the 00. The structure $\mathcal{C}_{i}=\lim_{s}\mathcal{C}_{i,s}$ clearly is as required.
@@
Of course the way Lemma 14 and Lemma 15 are stated we can not do this right away. We require more uniformity. However, looking at the proof of Lemma 14 the only step that is not uniform is the assumption that the first element is 00. However, in our case this is no obstacle, as we have defined $(\mathcal{C}_{i})_{i\in\omega}$ so that all $\mathcal{C}_{i}$ have 00 as its first element. ∎
@@
We use Corollary 17 to obtain computable $\mathsf{PA}$-verifiable linear order $\mathcal{L}_{1}\simeq\omega^{n}$ such that $\mathsf{PA}+\mathsf{TI}(\mathcal{L}_{1})\vdash\varphi$. Next we consider order $\mathcal{L}$ that is $\mathcal{L}_{1}$ with one additional element added on the very top, $\mathcal{L}\simeq\omega^{n}+1$. By Lemma 3 then we see that $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$ proves all instances of $\mathsf{TI}(\mathcal{L}_{1})$ and thus $\varphi$, which concludes the proof of Theorem 1.1. ∎

### v2

@@ 4. A theorem of Ash and Knight in $\mathsf{ACA_{0}}$
The goal of this section is to prove a version of a theorem of Ash and Knight [ash1990, Example 3] from computable structure theory in $\mathsf{ACA_{0}}$. In its original formulation, this theorem connects membership in hyperarithmetical sets to the well-orderedness of certain linear orders.
We are now ready to state the version of Ash and Knight’s theorem we have in mind. In the statement of this theorem and what follows $\eta$ denotes the order-type of the standard ordering on $\mathbb{Q}$.
@@
In order to prove Theorem 10 we will need a few more lemmas. The first is the relativized version of Shoenfield’s classical limit lemma. See [soare1999] for the classical proof of this result and more background on computability theory.
@@
A linear order $\mathcal{L}=(L,\preceq^{\mathcal{L}})$ is $\Delta^{0}_{n}(X)$ if $\preceq^{\mathcal{L}}$ is $\Delta^{0}_{n}(X)$. Notice that this implies that the universe $L$ is also $\Delta^{0}_{n}(X)$, since by reflexivity $x\in L$ iff $x\preceq^{\mathcal{L}}x$. If $\mathcal{L}$ is $\Delta^{0}_{1}(X)$, then we say that $\mathcal{L}$ is $X$-computable. We will denote linear orders by calligraphic letters and their universes $L$ by the corresponding capital letters. A sequence $(\mathcal{A}_{i})_{i\in\omega}$ of linear orders is uniformly $\Delta^{0}_{n}(X)$ if there is a $\Delta^{0}_{n}(X)$ set $S$ such that its $i^{\text{th}}$ column is equal to $\preceq^{\mathcal{A}_{i}}$, i.e.,
@@
From this, one can easily extract a pair of $\Pi^{0}_{2}$ formulas which, given parameter $X$, define the membership relation of $Y$, respectively, its negation. Thus $Y$ is $\Delta^{0}_{2}(X)$. ∎
From Lemma 11 we will now derive a limit lemma for linear orders.
Definition 12.
A sequence $(\mathcal{A}_{i})_{i\in\omega}$ of finite linear orders is locally coherent if $\preceq^{\mathcal{A}_{i}}\upharpoonright(A_{i}\cap A_{i+1})=\preceq^{\mathcal{A}_{i+1}}\upharpoonright(A_{i}\cap A_{i+1})$, for each $i$.
We will be interested in the situation when a linear order $\mathcal{L}$ is the limit of a locally coherent sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$. In this case, we think about $\mathcal{L}$ as the order obtained by the following procedure. We start with the order $\mathcal{A}_{0}$. At each stage $i$ we switch from $\mathcal{A}_{i}$ to $\mathcal{A}_{i+1}$ by removing all the elements $x\in A_{i}\setminus A_{i+1}$ and adding all the elements in $x\in A_{i+1}\setminus A_{i}$. The order $\mathcal{L}$ consists of all the elements $x$ that are present in some $\mathcal{A}_{i}$ and aren’t removed in any $\mathcal{A}_{j}$, for $j>i$. Whether $x\preceq^{\mathcal{L}}y$ can be checked in any $\mathcal{A}_{i}$ such that both $x,y$ aren’t removed at any stage $\geq i$.
@@
Note that for a linear order without successor relation $\mathcal{L}=\lim\mathcal{A}_{i}$ and successive $x,y\in\mathcal{L}$, new elements $z$ might appear and disappear between $x$ and $y$ inside $\mathcal{A}_{i}$’s for indefinitely large $i$. However, for a linear order with successor relation $\mathcal{L}=\lim\mathcal{A}_{i}$ and neighboring elements $x,y\in\mathcal{L}$, for large enough stages (after the point when both $x$ and $y$ will no longer be removed) we can guarantee that no more new elements will appear in $\mathcal{A}_{i}$’s between $x$ and $y$.
@@
Fix a set $X$. A linear order $\mathcal{L}$ is $\Delta^{0}_{2}(X)$ if and only if there is an $X$-computable locally coherent sequence $(\mathcal{A}_{i})_{i\in\omega}$ of finite linear orders with limit $\mathcal{L}$, i.e.,
$x\in L$ $\leftrightarrow\exists i(\forall j>i)\ x\in A_{j}$
$x\not\in L$ $\leftrightarrow\exists i(\forall j>i)\ x\not\in A_{j}$
$x\preceq^{\mathcal{L}}y$ $\leftrightarrow\exists i(\forall j>i)\ x\preceq^{\mathcal{A}_{j}}y.$
@@
()($\Rightarrow$)( ⇒ ). From Lemma 11 we get a sequence $(B_{i})_{i\in\omega}$ of finite sets with limit $\preceq^{\mathcal{L}}$. We may assume that no $B_{i}$ contains elements greater than $i$. We define a locally coherent sequence $(\mathcal{A}_{i})_{i\in\omega}$ as follows. Let $\mathcal{A}_{0}$ be the empty order. Assume we have defined $\mathcal{A}_{i}$. To define $\mathcal{A}_{i+1}$ first let $C_{i+1}$ be the largest subset of $\{x:\langle x,x\rangle\in B_{i+1}\}$ linearly ordered by $B_{i+1}$. Then let
$A_{i+1}=C_{i+1}\text{ and }\preceq^{\mathcal{A}_{i+1}}=B_{i+1}\upharpoonright A_{i+1}.$
This sequence is clearly uniformly $X$-computable. To see that it has limit $\mathcal{L}$, first consider $x,y\in L$ such that $x\preceq^{\mathcal{L}}y$. There exists $i$ such that for all $j>i$ and $z,w\leq\max(x,y)$ we have $\langle z,w\rangle\in B_{j}\iff z\preceq^{\mathcal{L}}w$. By definition, we then have that $x\preceq^{\mathcal{A}_{j}}y$. On the other hand, suppose that $x\not\in L$, then there is $i$ such that for all $j>i$, $\langle x,x\rangle\not\in B_{i}$. So for all $j>i$, $x\not\in A_{j}$.
@@
Fix a set $X$. A linear order $\mathcal{L}$ with successor relation $S^{\mathcal{L}}$ is $\Delta^{0}_{2}(X)$ if and only if there is an $X$-computable locally coherent sequence $(\mathcal{A}_{i})_{i\in\omega}$ of finite linear orders with successor relation with limit $\mathcal{L}$.
We are now ready to prove two lemmas crucial for our proof of Theorem 10. They give a sufficient condition for the existence of computable linear orderings of order-type $\omega\cdot\mathcal{L}$ for any linear ordering $\mathcal{L}$. Variations of these lemmas were first proven by Fellner [fellner1977].
@@
Fix $X$ and let $\mathcal{L}$ be a $\Delta^{0}_{2}(X)$ linear order with a least element. Then there is an $X$-computable linear order of order-type $\omega\cdot\mathcal{L}$ with an $X$-computable successor relation and 00 as its least element. [^5] We would like to thank Patrick Lutz and the anonymous referee who independently spotted a gap in the proof of Lemma 15 from an earlier version of the paper.
@@
Assume without loss of generality that 00 is the least element in $\mathcal{L}$. As $\mathcal{L}$ is $\Delta^{0}_{2}(X)$, by Corollary 13 there is a uniformly $X$-computable locally coherent sequence $(\mathcal{A}_{i})_{i\in\omega}$ of linear orders with limit $\mathcal{L}$. We may furthermore assume that $\mathcal{A}_{i}$ does not contain elements greater than $i$ and that each $\mathcal{A}_{i}$ has the correct information about 00.
@@
We define $\mathcal{L}^{\prime}$ of order-type $\omega\cdot\mathcal{L}$ as the union $\bigcup_{s\in\mathbb{N}}\mathcal{L}_{s}^{\prime}$ of an $X$-computable sequence of finite linear orders $\mathcal{L}_{0}^{\prime}\subseteq\mathcal{L}_{1}^{\prime}\subseteq\ldots$ where $\mathcal{L}_{i}^{\prime}$ has partial successor relation $S_{i}$. Together with $\mathcal{L}_{s}^{\prime}$ we define an $X$-computable sequence of order-preserving surjections $f_{s}\colon\mathcal{L}_{s}^{\prime}\to\mathcal{A}_{s}$ with the goal that $f=\lim f_{s}:\mathcal{L}^{\prime}\to\mathcal{L}$ is a surjective order-preserving function labeling the “$\omega$-blocks” of $\mathcal{L}^{\prime}$ with elements of $\mathcal{L}$.
We put $\mathcal{L}_{0}^{\prime}=\mathcal{A}_{0}$ and let $f_{0}$ be the identity function. We will now define $\mathcal{L}_{s+1}^{\prime}$ and $S_{s+1}$ as extensions of $\mathcal{L}_{s}^{\prime}$ and $S_{s}$ respectively and $f_{s+1}$ given $f_{s}$ as follows. During the construction, we will add fresh elements to $\mathcal{L}_{s+1}^{\prime}$ where a natural number $n$ is a fresh element if $n>s+1$ and has not been used in the construction before.
  1. (1)
    If there is a least natural number $a_{0}\in A_{s}$ such that the subordering of $\mathcal{A}_{s}$ consisting of natural numbers less than or equal to $a_{0}$ is not a subordering of $\mathcal{A}_{s+1}$, then let $\mathcal{A}^{\prime}$ be the suborder of $\mathcal{A}_{s+1}$ consisting of natural numbers less than $a_{0}$ and go to (2). Otherwise, let $\mathcal{A}^{\prime}=\mathcal{A}_{s}$ and go directly to (3).
  2. (2)
    For all elements $b\in A_{s}$ such that $b\geq a$ and all $x\in L_{s}^{\prime}$ with $f_{s}(x)=b$, let $f_{s+1}(x)$ be the next element to the left of $b$ in $\mathcal{A}^{\prime}$. If there are successive elements $x,y\in L_{s}^{\prime}$ with $f_{s+1}(x)=f_{s+1}(y)$ for which $S_{s}$ is undefined, set $S_{s+1}(x,y)$.
  3. (3)
    For each element $a\in A_{s+1}\setminus A^{\prime}$ we add to $\mathcal{L}_{s+1}^{\prime}$ a fresh element $x_{a}$ such that $f_{s+1}(x_{a})=a$. We compare $x_{a}$ with other elements $y\in\mathcal{L}_{s+1}^{\prime}$ as follows: $x_{a}\preceq^{\mathcal{L}^{\prime}_{s+1}}y$ iff $a\preceq^{A_{s+1}}f_{s+1}(y)$.
  4. (4)
    At last, grow every labeled block by an element to the right. I.e., for every $a\in L_{s+1}^{\prime}$ such that for all $b\succeq^{\mathcal{L}_{s+1}^{\prime}}a$, $f_{s+1}(b)\neq f_{s+1}(a)$, add a fresh element $x_{a}$, set $f_{s+1}(x_{a})=f_{s+1}(a)$, $S_{s+1}(a,x_{a})$ and define $\preceq^{\mathcal{L}_{s+1}^{\prime}}$ accordingly.
This finishes the constructions of $\mathcal{L}_{s+1}^{\prime}$, $S_{s+1}$ and $f_{s+1}$. Let $\mathcal{L}^{\prime}=\bigcup_{s\in\omega}\mathcal{L}_{s}^{\prime}$, $f=\lim f_{s}$ and $S=\bigcup_{s\in\omega}S_{s}$.
Verification: First, note that $\mathcal{L}^{\prime}$ is computable. For any two elements $x,y\in L^{\prime}$ with $m=\max\{x,y\}$, we have by construction that
$x\preceq^{\mathcal{L}^{\prime}}y\iff x\preceq^{\mathcal{L}^{\prime}_{m}}y.$
For the successor relation, the same reasoning applies except that at step (2) of the construction, we define the relation between elements added at previous stages. However, note that if at stage $s+1$, $S_{s+1}(x,y)$ is defined by step (2), then the element $x$ must be the rightmost element of its ”labeled block” and thus it was added by step (4) at stage $s$. Hence, for all $x,y\in L^{\prime}$ with $m=\max\{x,y\}$
$S(x,y)\iff S_{m+1}(x,y).$
Next, note that $f$ is well-defined. Whenever $f_{s+1}(x)\neq f_{s}(x)$, by construction $f_{s+1}(x)<f_{s}(x)$ and hence $f_{s}(x)$ stabilizes in the limit for any $x\in L^{\prime}$. To see that $range(f)=L$, first suppose that $a\in L$. By properties of $(\mathcal{A}_{i})_{i\in\omega}$, there is a least $i$ such that for all $j>i$, $\mathcal{A}_{i}\upharpoonright a=\mathcal{A}_{j}\upharpoonright a$. By construction $f_{i+1}(x)=a$ for some $x\in L_{i+1}$ and $f_{j}(x)=a$ for all $j>i+1$. On the other hand, suppose that $a\in range(f)\setminus L$. Then there is $i$ such that for all $j>i$, $\mathcal{A}_{i}\upharpoonright a=\mathcal{A}_{j}\upharpoonright a$. Thus, for all $j>i$, $a\not\in range(f_{j})$, a contradiction. That $f$ is order-preserving follows trivially from the construction. It remains to show that the interval of elements labeled with $a\in L$ is of order-type $\omega$. To see this let $i$ be such that there is $x\in L_{i}^{\prime}$ with $f_{j}(x)=a$ for all $j>i$. After stage $i$, the interval of elements with $f(x)=a$ will grow to the right by at least one and at most finitely many elements at each stage, but no elements will be added to the left or in between elements in the interval. Thus the order-type of each labeled interval is $\omega$. As $f$ is order-preserving and onto $\mathcal{L}$, it follows that $\mathcal{L}^{\prime}$ has order-type $\omega\cdot\mathcal{L}$ as required.
@@
Fix $X$ and let $\mathcal{L}$ be a $\Delta^{0}_{2}(X)$ linear order of order-type $\omega\cdot\mathcal{K}$ for some linear order $\mathcal{K}$. Further, assume that $\mathcal{L}$ has 00 as its least element and that the successor relation in $\mathcal{L}$ is $\Delta^{0}_{2}(X)$. Then there is an $X$-computable linear order isomorphic to $\mathcal{L}$ with 00 as its least element.
@@
As $\mathcal{L}$ is $\Delta^{0}_{2}(X)$, by Corollary 14 there is a uniformly $X$-computable locally coherent sequence of finite linear orders $(\mathcal{A}_{i})_{i\in\omega}$ with partial successor relation and $\mathcal{L}$ as its limit. Furthermore, we can assume without loss of generality that 00 is the least element of all $\mathcal{A}_{i}$.
@@
Verification: Clearly $\mathcal{L}^{\prime}$ is $X$-computable. It remains to show that $\mathcal{L}^{\prime}\cong\mathcal{L}$. First of all notice that the order $\mathcal{L}^{\prime}$ splits into the blocks $B_{a}$, for $a\in\mathcal{L}$, where $B_{a}=\bigcup\limits_{s\geq s_{a}}f_{s}^{-1}(a)$ and $s_{a}$ is the first stage such that $a\in\bigcap\limits_{s\geq s_{a}}A_{s}$. In order to finish the proof it is enough to show that all $B_{a}$ are finite and non-empty. For a given $a\in\mathcal{L}$ consider its successor $b$. In all the orders $\mathcal{A}_{s}$, for $s\geq\max(s_{a},s_{b})$, the element $b$ is the successor of $a$ according to $S^{\mathcal{A}_{s}}$. Thus no elements will appear in between $a$ and $b$ in the orders $\mathcal{A}_{s}$. Therefore, no new elements will be added to $B_{a}$ after the stage $\max(s_{a},s_{b})$, and hence $B_{a}$ is finite. The block $B_{a}$ is non-empty since by construction $f^{-1}_{s_{a}}(a)=\{x_{a}\}$.∎
Combining Lemmas 15 and 16 we get a formal version of a special case of a theorem due to Ash [ash1991] as stated in [ash2000, Theorem 9.11]: If a linear order $\mathcal{L}$ has an isomorphic copy that is $\Delta^{0}_{3}$, then $\omega\cdot\mathcal{L}$ has a computable copy. The proof of this in two steps, first applying Lemma 15 with $X=\emptyset^{\prime}$ and then using Lemma 16, was communicated to us by Andrey Frolov.
@@
We can now build a uniformly $\Delta_{2n}$ sequence of linear orders $(\mathcal{C}_{i})_{i\in\omega}$ such that if $x\in X$, then $\mathcal{C}_{i}$ is the linear order only containing the element 00 and if $x\not\in X$, then $\mathcal{C}_{i}$ has order-type $1+\eta$ where the first element is 00. The construction is again in stages where $\mathcal{C}_{i,s}$ contains only one element 00 if for all $y<s$ $\langle x,y\rangle\in Y$ and otherwise we build a copy of $\eta$ after the 00. The structure $\mathcal{C}_{i}=\lim_{s}\mathcal{C}_{i,s}$ clearly is as required.
@@
Of course, the way Lemma 15 and Lemma 16 are stated, we can not do this right away. We require more uniformity. However, looking at the proof of Lemma 15 the only step that is not uniform is the assumption that the first element is 00. However, in our case, this is no obstacle, as we have defined $(\mathcal{C}_{i})_{i\in\omega}$ so that all $\mathcal{C}_{i}$ have 00 as its first element. ∎
@@
We use Corollary 18 to obtain computable $\mathsf{PA}$-verifiable linear order $\mathcal{L}_{1}\simeq\omega^{n}$ such that $\mathsf{PA}+\mathsf{TI}(\mathcal{L}_{1})\vdash\varphi$. Next, we consider the order $\mathcal{L}$ obtained from $\mathcal{L}_{1}$ by adding a new greatest element, i.e., $\mathcal{L}\simeq\omega^{n}+1$. By Lemma 3 we see that $\mathsf{RFN}^{\mathcal{L}}(\mathsf{PA})$ proves all instances of $\mathsf{TI}(\mathcal{L}_{1})$ and thus $\varphi$, which concludes the proof of Theorem 1.1. ∎